where does square root 8 go on a number line.
Found 2 solutions by checkley71, invisiblebug590:
Answer by checkley71(8403)

SQRT8=SQRT(4*2)=2SQRT2=2*1.414=2.828 ANSWER.

Answer by invisiblebug590(1)

some where near 2.828